# State-space models are accurate and efficient neural operators for dynamical systems

### v1

@@ 4.8.3 Learning the PK-PD Model under Limited Data
Due to the constraints of real-world biomedical experiments, the availability of PK-PD model data is often limited. The high cost of conducting experiments, the ethical considerations involved, and the restrictions on data sharing contribute to this scarcity, making it essential to develop models that can efficiently utilize limited data. To this end, we conduct an additional study on Mamba to investigate its generalization with respect to training dataset size under a fixed test set. We generate 5000 data in total as in Section 4.8.2 and leave 1000 as the fixed test set. We manipulate the train set size within 4000, 2000, 1000, 500, 250, and 125. Regarding Mamba, we opt for a block with skip connection and layer normalization, with 16 model dimensions and 16 intermediate dimensions, which have 9697 parameters. The results shown in Table 10 demonstrate that Mamba’s test error and generalization gap become larger with decreasing training data while the training loss is decreasing. As we have less training data, fitting and overfitting the training set becomes easier for the Mamba model, leading to decreasing training loss and increasing the generalization gap. However, it can still achieve some good results under limited data settings.
Train Data Num 4000 2000 1000 500 250 125 125
Mamba Params Num 9697 9697 9697 9697 9697 9697 8865
Train Rel. $L_{2}$ Error 1.605E-3 1.731E-3 1.564E-3 1.436E-3 1.132E-3 7.448E-4 1.975E-3
Test Rel. $L_{2}$ Error 1.685E-3 2.151E-3 3.041E-3 7.132E-3 1.228E-2 2.158E-2 1.685E-2
Table 10: Results for the PK-PD model under limited data setting corresponding to Section 4.8.3. The test error drops in the last column since we reduced the model size to prevent overfitting.
To mitigate the generalization gap under data-limited settings, we can reduce the model size and parameter number to mitigate the overfitting effect. Hence, we train another smaller Mamba model with two blocks and only 8 model dimensions and 8 intermediate dimensions, which have 8865 parameters, for the 125 training data setting. The smaller model achieves a larger training error but a lower testing error and, thus, a smaller generalization gap. Smaller models exhibit weaker fitting capabilities, as evidenced by larger training losses; however, they are also less prone to overfitting. This highlights the trade-off between model overfitting and expressivenes. In practical applications, a portion of the training set can be separated as a validation set to determine the appropriate model size.
@@ 4.8.4 Extrapolation to treatment schedule parameters
In this study, we performed an extrapolation of various treatment schedule parameters to evaluate the efficacy of different dosing regimens. The analysis included three distinct datasets, each containing 200 out-of-distribution test data, and was designed to test the model’s robustness and predictive capability under various conditions. The administration dose was maintained within the therapeutic range to avoid the toxic zone, so we did not change this range. Data was generated to ensure that all dose amounts were uniformly distributed across the first and second datasets, differing from the training dataset, where doses were selected from a random uniform function. The input values for the treatment plan were modified to study the efficacy of different schedules in specific populations. Since the population remains the same, the PD parameters were kept within the training range using random uniform values. The first dataset includes treatment schedules with intervals that are out-of-training-range (e.g., every 7 and 8 days) and first administration days that are out-of-training-range (e.g., day 14, 15, or 16). Within the training time range of [0, 35], all patients receive only two doses in a multi-dose administration. In the second dataset, the input values for the first day of administration were set out of range, while the schedule for the second dataset remained the same (e.g., day 2 to day 6). This approach was used to test the impact of starting cancer treatment later than in the training dataset. In the third dataset, the dose is chosen from a random uniform function similar to the training dataset. The same schedule is applied but with a different first day of administration to ensure all patients receive three treatments within the [0, 35] day range. In addition, the training set is the 5,000 data generated in Section 4.8.2, which is a different distribution from all three test sets.
@@ 4.8.5 Physics-Informed Mamba Operator Learning
Motivation and Experimental Setting. As we mentioned, the PK-PD model data require real-world laboratory experiments; we need to learn the PK model under limited data. To investigate the efficacy of training the Mamba model with very limited data, specifically with only 5 labeled training samples and 50 test samples, we explored how incorporating physical information from the ODE might enhance model performance. In addition to the labeled data, we utilized 45 unlabeled data points, which provide input information without corresponding labels. All data are generated according to Section 4.8.1. Specifically, the loss function we employed consists of a data loss term for the labeled data and a physics loss term for the unlabeled data. The time-derivative in the physics-informed loss is computed via the central finite difference scheme. The data and physics-informed losses are all given unity weights. Our objective is to examine whether integrating physical information during model training can improve Mamba’s generalization capability and assess its practical applicability in real-world scenarios.
Results. We obtain a relative $L_{2}$ error of 8.718E-2 with data loss only. If the physics-informed loss on the unlabeled data is employed in addition to the data loss, i.e., the hybrid method with both data loss and physics loss, we improve the error to 1.872E-2. However, if we discard the labeled data loss and only train on the physics-informed loss on the unlabeled data, then we obtain an error of 8.791E-1. In the physics-informed loss function, the ODE on $A_{1}(t)$ is stiff and thus rigid to learn due to the discontinuous $\mu$, which may not be an excellent supervision to guide Mamba to the ODE solution. In contrast, the data loss in the hybrid approach pins the model to the correct solution, making it the strongest training methodology. In conclusion, Mamba can be flexibly trained in data-driven, physics-informed, or hybrid scenarios. The hybrid approach is the best, thanks to blending data with physics. This demonstrates Mamba’s potential in handling real-world PK-PD models when experimental data is scarce by using physics to improve its data efficiency and generalization.
@@ 5 Summary and Discussion
Our findings underscore the potential of Mamba as a versatile and efficient tool for dynamical system modeling. By bridging the gap between traditional numerical solvers and modern machine learning techniques, Mamba opens new avenues for advancing scientific research and practical applications across multiple domains. Future work will explore further enhancements to the Mamba architecture, expanding its applicability and refining its performance on even more challenging dynamical systems and going beyond ODEs to PDEs. This can be readily accomplished by combining DeepOnet with Mamba or a Unet architecture with Mamba, e.g. following the work in [37], where the spatial representation is based on a Unet while conditioning on time using an extra MLP.
@@
Beyond the superior performance of Mamba on this plurality of benchmarks, we also demonstrated its effectiveness on an important real-world application. One of the most significant challenges in the field of Quantitative Systems Pharmacology is the prediction of parameters for known compartmental and non-compartmental models. This is particularly the case for parameters of PK and PD of patient-specific models for individuals within a population. The objective of our future work is to integrate the Mamba model as a predictive tool for parameter inference, with the aim of providing tailored results for each individual. Such integration would facilitate the assessment of disparate treatment regimen outcomes for each patient based on a limited set of patient-specific features. It is important to note that in the present study, a single PK model was used for all patients under the assumption that a specific drug was used. Traditionally, a single PK model is developed for each drug within a given population. However, in reality, the concentration of the same drug in plasma varies among individuals due to differences in ADME (absorption, distribution, metabolism, and excretion) properties of the same drug for each patient. With sufficient data, a more accurate and individualized model could be developed that better reflects real-world applications.
@@
In multi-dose regimens, resistance to chemotherapeutic agents such as paclitaxel is a common challenge. This phenomenon, known as chemotherapy resistance, occurs when tumors that initially respond to treatment begin to grow again. Resistance can result from pharmacodynamic resistance or pharmacokinetic tolerance, particularly in the context of multi-dose regimens. The complexity of modeling these adaptive mechanisms in PK and PD makes it difficult to accurately measure their impact on treatment efficacy. Because they either reduce the effective concentration of the drug or diminish the response of cancer cells to therapy, incorporating additional variables—such as the specific cancer type or other relevant clinical data—can improve the accuracy of patient-specific predictive models. Therefore, the Mamba models with linear cost and superior accuracy as robustness could be utilized effectively in such cases.
@@ A.1.3 Laplace Neural Operator (LNO)
where $\gamma_{n}=\beta_{n}\sum_{\ell=-\infty}^{\infty}\frac{\alpha_{\ell}}{\mu_{n}-i\omega_{\ell}}$ and $\lambda_{\ell}=\alpha_{\ell}\sum_{n=1}^{N}\frac{\beta_{n}}{i\omega_{\ell}-\mu_{n}}.$ Consequently, the output function in the original space can be computed via an inverse Laplace transform, and be represented by
@@
Model 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
GRU 5.726E-4 2.803E-3 1.064E-2 1.376E-2 1.522E-2 2.172E-2 1.970E-2 2.897E-2 1.827E-2 2.364E-2
LSTM 8.731E-4 4.536E-3 1.224E-2 1.697E-2 1.946E-2 2.868E-2 2.446E-2 3.446E-2 2.395E-2 2.898E-2
DeepONet 2.602E-3 9.120E-3 1.827E-2 2.668E-2 3.141E-2 4.326E-2 3.550E-2 4.979E-2 3.341E-2 3.898E-2
FNO 5.170E-2 5.082E-2 6.421E-2 8.281E-2 8.123E-2 9.682E-2 1.174E-1 1.475E-1 1.524E-1 1.596E-1
LNO 2.262E-3 1.335E-2 3.277E-2 5.004E-2 5.748E-2 8.480E-2 6.853E-2 9.851E-2 6.502E-2 7.874E-2
Transformer 1.471E-2 2.102E-2 4.109E-2 5.312E-2 5.853E-2 7.835E-2 7.078E-2 9.047E-2 6.731E-2 8.003E-2
Oformer-V 6.858E-4 7.029E-4 2.542E-3 2.953E-3 2.953E-3 5.286E-3 5.536E-3 7.817E-3 3.816E-3 6.808E-3
Oformer-G 6.460E-4 4.137E-4 7.709E-4 8.221E-4 8.444E-4 1.347E-3 1.415E-3 1.549E-3 1.065E-3 1.729E-3
Oformer-F 5.573E-4 2.949E-4 4.263E-4 4.767E-4 5.843E-4 8.541E-4 7.996E-4 2.762E-3 7.379E-4 9.089E-4
GNOT 1.028E-3 8.031E-4 2.186E-3 1.879E-3 2.316E-3 5.353E-3 6.543E-3 7.115E-3 2.936E-3 8.543E-3
Mamba 2.761E-4 2.982E-4 3.517E-4 3.830E-4 4.017E-4 4.370E-4 4.138E-4 4.661E-4 3.993E-4 4.103E-4
@@
Model 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
GRU 8.735E-2 3.895E-2 2.240E-2 8.818E-3 6.150E-3 3.097E-3 1.255E-3 8.555E-4 3.970E-4 4.207E-4
LSTM 5.736E-2 2.387E-2 1.268E-2 6.553E-3 3.963E-3 1.806E-3 1.059E-3 9.732E-4 4.137E-4 4.098E-4
DeepONet 2.275E-1 7.527E-2 2.542E-2 9.500E-3 4.099E-3 1.893E-3 9.797E-4 3.647E-3 3.303E-4 1.028E-3
FNO 8.780E-1 4.551E-1 2.248E-1 1.088E-1 5.593E-2 3.371E-2 3.236E-2 2.045E-2 2.555E-2 1.734E-2
LNO 1.287E-1 4.425E-2 1.706E-2 7.268E-3 3.624E-3 1.905E-3 1.085E-3 1.421E-3 4.383E-4 5.512E-4
Transformer 6.722E-1 3.210E-1 1.589E-1 6.367E-2 2.669E-2 1.204E-2 7.776E-3 6.345E-3 5.544E-3 5.325E-3
Oformer-V 1.050E-1 3.761E-2 1.449E-2 5.971E-3 3.060E-3 1.401E-3 7.682E-4 5.621E-4 3.209E-4 2.672E-4
Oformer-G 6.097E-2 2.226E-2 9.664E-3 3.947E-3 2.300E-3 9.772E-4 5.050E-4 3.272E-4 2.101E-4 1.738E-4
Oformer-F 8.852E-2 3.096E-2 1.494E-2 5.835E-3 4.121E-3 1.955E-3 6.925E-4 3.597E-4 1.891E-4 1.709E-4
GNOT 1.282E-1 5.437E-2 2.664E-2 1.227E-2 6.648E-3 2.580E-3 1.139E-3 7.249E-4 3.514E-4 2.967E-4
Mamba 2.002E-2 6.908E-3 3.260E-3 1.522E-3 8.893E-4 4.889E-4 3.024E-4 2.226E-4 1.510E-4 1.369E-4
@@
Model 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
GRU 6.960E-4 5.940E-3 1.866E-2 1.951E-2 2.393E-2 2.662E-2 3.037E-2 3.911E-2 3.857E-2 3.665E-2
LSTM 1.226E-3 1.634E-3 6.682E-4 3.305E-3 4.683E-3 5.804E-3 2.783E-3 7.751E-3 9.388E-3 6.308E-3
DeepONet 1.164E-3 7.931E-3 1.701E-2 1.986E-2 2.246E-2 2.736E-2 2.843E-2 3.656E-2 3.571E-2 3.253E-2
FNO 6.648E-2 6.030E-2 6.648E-2 7.215E-2 7.949E-2 8.679E-2 9.242E-2 1.061E-1 1.042E-1 1.209E-1
LNO 1.458E-3 9.632E-3 2.612E-2 3.357E-2 4.058E-2 4.967E-2 4.971E-2 6.880E-2 5.947E-2 5.947E-2
Transformer 1.787E-2 2.193E-2 4.227E-2 4.723E-2 5.607E-2 6.306E-2 6.927E-2 8.232E-2 8.198E-2 7.737E-2
Oformer-V 4.073E-4 4.902E-4 9.225E-4 1.540E-3 1.855E-3 2.399E-3 2.101E-3 3.270E-3 3.605E-3 3.006E-3
Oformer-G 4.086E-4 5.932E-4 6.447E-4 6.558E-4 7.338E-4 7.490E-4 7.522E-4 8.885E-4 8.962E-4 8.121E-4
Oformer-F 4.063E-4 4.296E-4 4.188E-4 4.022E-4 4.430E-4 4.223E-4 4.262E-4 4.835E-4 4.765E-4 4.621E-4
GNOT 2.061E-3 1.069E-3 1.395E-3 1.566E-3 1.919E-3 2.134E-3 1.934E-3 2.638E-3 3.099E-3 2.529E-3
Mamba 1.204E-4 1.226E-4 1.431E-4 1.435E-4 2.589E-4 1.690E-4 1.829E-4 2.077E-4 2.409E-4 2.188E-4

### v2

@@ 4.8.3 Learning the PK-PD Model under Limited Data
Due to the constraints of real-world biomedical experiments, the availability of PK-PD model data is often limited. The high cost of conducting experiments, the ethical considerations involved, and the restrictions on data sharing contribute to this scarcity, making it essential to develop models that can efficiently utilize limited data. To this end, we conduct an additional study on Mamba to investigate its generalization with respect to training dataset size under a fixed test set. We generate 5000 data in total as in Section 4.8.2 and leave 1000 as the fixed test set. We manipulate the train set size within 4000, 2000, 1000, 500, 250, and 125. Regarding Mamba, we opt for a block with skip connection and layer normalization, with 16 model dimensions and 16 intermediate dimensions, which have 9697 parameters, as in the previous subsection. We keep all the training settings the same as in the previous subsection except for the number of training epochs. We maintain the product between train set size and the number of training epochs to be 4000 $\times$ 10,000 to maintain the total iterations similar across various settings given less data. For instance, the training epochs number for 2000 data is 20,000. The results shown in Table 11 demonstrate that Mamba’s test error and generalization gap become larger with decreasing training data while the training loss is decreasing. As we have less training data, fitting and overfitting the training set becomes easier for the Mamba model, leading to decreasing training loss and increasing the generalization gap. However, it can still achieve some good results under limited data settings.
Train Data Num 4000 2000 1000
Mamba Params Num 9697 9697 9697
Train Rel. $L_{2}$ Error 1.605E-3$\pm$2.514E-4 1.512E-3$\pm$1.651E-4 1.289E-3$\pm$2.417E-4
Test Rel. $L_{2}$ Error 1.685E-3$\pm$2.929E-4 1.940E-3$\pm$2.243E-4 2.873E-3$\pm$1.575E-4
Train Data Num 500 250 125
Mamba Params Num 9697 9697 9697
Train Rel. $L_{2}$ Error 1.129E-3$\pm$2.668E-4 1.027E-3$\pm$1.479E-4 7.137E-4$\pm$1.038E-4
Test Rel. $L_{2}$ Error 6.633E-3$\pm$5.850E-4 1.444E-2$\pm$3.052E-3 2.009E-2$\pm$1.945E-3
Table 11: Results for the PK-PD model under limited data setting corresponding to Section 4.8.3.
@@ 4.8.4 Extrapolation to treatment schedule parameters
In this study, we performed an extrapolation of various treatment schedule parameters to evaluate the efficacy of different dosing regimens. The analysis included three distinct datasets, each containing 200 out-of-distribution test data, and was designed to test the model’s robustness and predictive capability under various conditions. The administration dose was maintained within the therapeutic range to avoid the toxic zone, so we did not change this range. Data was generated to ensure that all dose amounts were uniformly distributed across the first and second datasets, differing from the training dataset, where doses were selected from a random uniform function. The input values for the treatment plan were modified to study the efficacy of different schedules in specific populations. Since the population remains the same, the PD parameters were kept within the training range using random uniform values. The first dataset includes treatment schedules with intervals that are out-of-training-range (e.g., every 7 and 8 days) and first administration days that are out-of-training-range (e.g., day 14, 15, or 16). Within the training time range of [0, 35], all patients receive only two doses in a multi-dose administration. In the second dataset, the input values for the first day of administration were set out of range, while the schedule for the second dataset remained the same (e.g., day 2 to day 6). This approach was used to test the impact of starting cancer treatment later than in the training dataset. In the third dataset, the dose is chosen from a random uniform function similar to the training dataset. The same schedule is applied but with a different first day of administration to ensure all patients receive three treatments within the [0, 35] day range. In addition, the training set is the 5,000 data generated in Section 4.8.2, which is a different distribution from all three test sets. We choose the top three models in the IID test in Section 4.8.2: GRU, LSTM, and Mamba. We also adopt the same model architecture and optimization procedure as in Section 4.8.2.
@@ 4.8.5 Physics-Informed Mamba Operator Learning
Motivation and Experimental Setting. As we mentioned, the PK-PD model data require real-world laboratory experiments; we need to learn the PK model under limited data. To investigate the efficacy of training the Mamba model with very limited data, specifically with only five labeled training samples and 50 test samples, we explored how incorporating physical information from the ODE might enhance model performance. In addition to the labeled data, we utilized 45 unlabeled data points, which provide input information without corresponding labels. All data are generated according to Section 4.8.1. Specifically, the loss function we employed consists of a data loss term for the labeled data and a physics loss term for the unlabeled data. The time-derivative in the physics-informed loss is computed via the central finite difference scheme. The data and physics-informed losses are all given unity weights. Our objective is to examine whether integrating physical information during model training can improve Mamba’s generalization capability and assess its practical applicability in real-world scenarios. The Mamba structure and optimization procedure are the same as Section 4.8.2 except the training epochs number is prolonged to 100,000 due to the smaller dataset.
Results. We obtain a relative $L_{2}$ error of 8.718E-2$\pm$2.591E-2 with data loss only. If the physics-informed loss on the unlabeled data is employed in addition to the data loss, i.e., the hybrid method with both data loss and physics loss, we improve the error to 1.872E-2$\pm$1.023E-2. However, if we discard the labeled data loss and only train on the physics-informed loss on the unlabeled data, then we obtain an error of 8.791E-1$\pm$1.576E-1. In the physics-informed loss function, the ODE on $A_{1}(t)$ is stiff and thus rigid to learn due to the discontinuous $\mu$, which may not be an excellent supervision to guide Mamba to the ODE solution. In contrast, the data loss in the hybrid approach pins the model to the correct solution, making it the strongest training methodology. In conclusion, Mamba can be flexibly trained in data-driven, physics-informed, or hybrid scenarios. The hybrid approach is the best, thanks to blending data with physics. This demonstrates Mamba’s potential in handling real-world PK-PD models when experimental data is scarce by using physics to improve its data efficiency and generalization.
@@ 5 Summary and Discussion
Our findings underscore the potential of Mamba as a versatile and efficient tool for dynamical system modeling. By bridging the gap between traditional numerical solvers and modern machine learning techniques, Mamba opens new avenues for advancing scientific research and practical applications across multiple domains. Future work will explore further enhancements to the Mamba architecture, expanding its applicability and refining its performance on even more challenging dynamical systems and going beyond ODEs to PDEs. This can be readily accomplished by combining DeepOnet with Mamba or a Unet architecture with Mamba, e.g., following the work in [38], where the spatial representation is based on a Unet while conditioning on time using an extra MLP.
@@
Beyond the superior performance of Mamba on this plurality of benchmarks, we also demonstrated its effectiveness in an important real-world application. One of the most significant challenges in the field of Quantitative Systems Pharmacology is the prediction of parameters for known compartmental and non-compartmental models. This is particularly the case for parameters of PK and PD of patient-specific models for individuals within a population. The objective of our future work is to integrate the Mamba model as a predictive tool for parameter inference, whether constant or time-varying [9], with the aim of providing tailored results for each individual. Such integration would facilitate the assessment of disparate treatment regimen outcomes for each patient based on a limited set of patient-specific features. It is important to note that in the present study, a single PK model was used for all patients under the assumption that a specific drug was used. Traditionally, a single PK model is developed for each drug within a given population. However, in reality, the concentration of the same drug in plasma varies among individuals due to differences in ADME (absorption, distribution, metabolism, and excretion) properties of the same drug for each patient. With sufficient data, a more accurate and individualized model could be developed that better reflects real-world applications.
@@
In multi-dose regimens, resistance to chemotherapeutic agents such as paclitaxel is a common challenge. This phenomenon, known as chemotherapy resistance, occurs when tumors that initially respond to treatment begin to grow again. Resistance can result from pharmacodynamic resistance or pharmacokinetic tolerance, particularly in the context of multi-dose regimens. The complexity of modeling these adaptive mechanisms in PK and PD makes it difficult to accurately measure their impact on treatment efficacy. Because they either reduce the effective concentration of the drug or diminish the response of cancer cells to therapy, incorporating additional variables—such as the specific cancer type or other relevant clinical data—can improve the accuracy of patient-specific predictive models. Therefore, the Mamba models with linear cost, superior accuracy, and robustness could be utilized effectively in such cases.
@@ A.1.3 Laplace Neural Operator (LNO)
where $\gamma_{n}=\beta_{n}\sum_{\ell=-\infty}^{\infty}\frac{\alpha_{\ell}}{\mu_{n}-i\omega_{\ell}}$ and $\lambda_{\ell}=\alpha_{\ell}\sum_{n=1}^{N}\frac{\beta_{n}}{i\omega_{\ell}-\mu_{n}}.$ Consequently, the output function in the original space can be computed via an inverse Laplace transform and be represented by
@@
Model Metric 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
GRU Mean 5.992E-04 2.494E-03 1.112E-02 1.442E-02 1.628E-02 2.401E-02 2.107E-02 3.049E-02 2.010E-02 2.581E-02
Std 2.301E-05 3.938E-04 1.322E-03 2.017E-03 1.990E-03 2.968E-03 1.595E-03 1.343E-03 2.435E-03 2.241E-03
LSTM Mean 7.893E-04 4.149E-03 1.225E-02 1.669E-02 1.930E-02 2.835E-02 2.432E-02 3.429E-02 2.393E-02 2.896E-02
Std 7.518E-05 3.347E-04 2.760E-04 4.666E-04 3.515E-04 5.668E-04 2.441E-04 2.258E-04 7.072E-04 4.910E-04
DeepONet Mean 1.949E-03 6.438E-03 1.325E-02 1.906E-02 2.253E-02 3.114E-02 2.541E-02 3.658E-02 2.361E-02 2.833E-02
Std 6.547E-04 2.976E-03 5.913E-03 8.861E-03 1.008E-02 1.392E-02 1.148E-02 1.515E-02 1.105E-02 1.234E-02
FNO Mean 6.335E-02 6.755E-02 8.991E-02 1.127E-01 1.160E-01 1.322E-01 1.536E-01 1.941E-01 1.980E-01 2.059E-01
Std 1.015E-02 1.470E-02 2.380E-02 3.336E-02 3.790E-02 3.641E-02 4.457E-02 5.259E-02 5.545E-02 5.798E-02
LNO Mean 1.942E-03 1.062E-02 2.728E-02 4.087E-02 4.838E-02 7.193E-02 5.745E-02 8.510E-02 5.308E-02 6.668E-02
Std 3.373E-04 2.681E-03 7.010E-03 1.093E-02 1.107E-02 1.617E-02 1.313E-02 1.582E-02 1.409E-02 1.459E-02
Transformer Mean 3.043E-02 2.553E-02 2.609E-02 3.160E-02 3.452E-02 4.393E-02 4.235E-02 5.624E-02 4.157E-02 4.598E-02
Std 9.951E-04 3.193E-03 3.195E-03 1.139E-02 1.291E-02 2.369E-02 2.077E-02 3.032E-02 2.073E-02 2.361E-02
Oformer-V Mean 6.149E-04 1.002E-03 2.460E-03 3.088E-03 3.985E-03 6.035E-03 5.370E-03 8.776E-03 4.563E-03 6.700E-03
Std 7.592E-05 3.519E-04 1.225E-03 1.687E-03 2.287E-03 3.649E-03 3.169E-03 5.271E-03 2.822E-03 4.020E-03
Oformer-G Mean 6.577E-04 3.970E-04 5.075E-04 5.177E-04 5.665E-04 7.167E-04 6.968E-04 2.258E-03 6.199E-04 8.159E-04
Std 1.853E-04 1.742E-04 1.748E-04 1.879E-04 1.409E-04 1.995E-04 2.986E-04 1.066E-03 2.723E-04 2.525E-04
Oformer-F Mean 6.134E-04 3.319E-04 3.853E-04 4.168E-04 4.133E-04 4.555E-04 4.138E-04 1.436E-03 3.660E-04 5.189E-04
Std 2.263E-04 1.528E-04 1.996E-04 2.517E-04 2.155E-04 2.440E-04 2.138E-04 1.067E-03 1.778E-04 3.079E-04
GNOT Mean 6.422E-04 5.473E-04 1.144E-03 1.345E-03 1.547E-03 2.516E-03 2.490E-03 4.184E-03 2.182E-03 2.978E-03
Std 1.465E-05 1.373E-04 4.592E-04 6.569E-04 7.168E-04 1.159E-03 1.105E-03 2.033E-03 9.995E-04 1.295E-03
Mamba Mean 2.776E-04 2.948E-04 3.412E-04 3.765E-04 3.921E-04 4.134E-04 3.907E-04 4.337E-04 3.709E-04 3.755E-04
Std 5.312E-06 4.985E-06 1.396E-05 2.851E-05 3.648E-05 5.663E-05 5.880E-05 8.012E-05 6.632E-05 7.554E-05
@@
Model Metric 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
GRU Mean 7.693E-02 3.499E-02 1.961E-02 7.521E-03 5.027E-03 2.408E-03 1.188E-03 8.439E-04 4.031E-04 4.249E-04
Std 7.460E-03 3.892E-03 3.346E-03 1.011E-03 9.871E-04 5.223E-04 1.868E-04 1.389E-04 2.473E-05 1.580E-05
LSTM Mean 5.494E-02 2.532E-02 1.421E-02 6.832E-03 4.327E-03 1.973E-03 1.146E-03 1.253E-03 4.167E-04 4.376E-04
Std 3.538E-03 3.071E-03 2.270E-03 9.035E-04 7.105E-04 2.483E-04 8.139E-05 2.041E-04 3.837E-06 5.258E-05
DeepONet Mean 2.333E-01 7.821E-02 2.864E-02 1.054E-02 4.455E-03 1.976E-03 9.950E-04 3.164E-03 3.709E-04 1.033E-03
Std 7.506E-02 2.721E-02 9.835E-03 2.740E-03 7.972E-04 2.257E-04 6.357E-05 1.109E-03 3.505E-05 2.519E-04
FNO Mean 8.276E-01 4.336E-01 2.942E-01 1.229E-01 3.033E-01 1.846E-01 3.387E-02 2.071E-02 2.557E-02 1.835E-02
Std 5.038E-02 2.147E-02 6.943E-02 1.413E-02 2.474E-01 1.509E-01 1.509E-03 2.601E-04 1.353E-05 1.010E-03
LNO Mean 1.298E-01 4.380E-02 1.648E-02 6.948E-03 3.417E-03 1.810E-03 1.048E-03 1.327E-03 4.259E-04 5.044E-04
Std 1.567E-03 6.290E-04 4.948E-04 2.661E-04 1.560E-04 7.509E-05 3.838E-05 1.576E-04 1.877E-05 3.626E-05
Transformer Mean 3.623E-01 2.312E-01 1.684E-01 1.134E-01 7.871E-02 4.260E-02 2.768E-02 1.758E-02 1.538E-02 1.473E-02
Std 2.580E-02 2.063E-02 1.765E-02 1.890E-02 9.946E-03 6.458E-03 3.333E-03 7.841E-04 4.882E-04 6.280E-04
Oformer-V Mean 3.562E-01 1.013E-01 3.438E-02 1.335E-02 6.245E-03 3.040E-03 1.569E-03 1.076E-03 5.498E-04 4.338E-04
Std 8.122E-02 1.176E-02 3.195E-03 1.250E-03 6.783E-04 3.392E-04 2.082E-04 9.859E-05 6.452E-05 5.031E-05
Oformer-G Mean 6.109E-02 2.411E-02 9.964E-03 4.436E-03 2.195E-03 1.102E-03 6.137E-04 4.572E-04 2.441E-04 2.179E-04
Std 6.217E-03 7.248E-04 4.695E-04 3.873E-04 2.692E-04 1.396E-04 9.158E-05 8.254E-05 4.565E-05 6.217E-05
Oformer-F Mean 5.853E-02 2.641E-02 1.183E-02 5.376E-03 2.648E-03 1.287E-03 7.096E-04 5.042E-04 2.708E-04 2.606E-04
Std 1.305E-03 7.313E-04 1.203E-03 1.033E-03 5.552E-04 2.691E-04 1.397E-04 1.110E-04 5.843E-05 6.325E-05
GNOT Mean 1.023E-01 3.958E-02 1.585E-02 6.551E-03 3.398E-03 1.509E-03 7.431E-04 5.171E-04 2.677E-04 2.241E-04
Std 1.641E-02 8.068E-03 3.411E-03 1.165E-03 5.416E-04 2.238E-04 7.435E-05 3.796E-05 1.373E-05 1.238E-05
Mamba Mean 1.758E-02 6.081E-03 2.699E-03 1.243E-03 7.350E-04 4.146E-04 2.679E-04 2.157E-04 1.466E-04 1.332E-04
Std 6.335E-03 1.933E-03 7.687E-04 3.262E-04 1.925E-04 9.272E-05 5.148E-05 3.555E-05 1.939E-05 1.467E-05
@@
Model Metric 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
GRU Mean 5.145E-04 4.388E-03 1.361E-02 1.547E-02 1.918E-02 2.193E-02 2.354E-02 3.185E-02 3.104E-02 2.881E-02
Std 1.675E-04 1.209E-03 4.441E-03 3.018E-03 3.722E-03 3.392E-03 5.762E-03 5.922E-03 6.153E-03 6.761E-03
LSTM Mean 7.893E-04 7.746E-03 1.517E-02 1.870E-02 2.280E-02 2.637E-02 2.686E-02 3.500E-02 3.466E-02 3.178E-02
Std 3.590E-04 5.503E-03 1.076E-02 1.115E-02 1.295E-02 1.461E-02 1.708E-02 1.931E-02 1.795E-02 1.814E-02
DeepONet Mean 1.135E-03 6.607E-03 1.427E-02 1.668E-02 1.885E-02 2.288E-02 2.379E-02 3.080E-02 3.008E-02 2.737E-02
Std 6.133E-05 9.810E-04 2.086E-03 2.273E-03 2.600E-03 3.181E-03 3.392E-03 4.174E-03 4.074E-03 3.787E-03
FNO Mean 5.158E-02 4.985E-02 5.726E-02 6.365E-02 7.224E-02 7.919E-02 8.325E-02 1.008E-01 9.730E-02 1.170E-01
Std 8.791E-04 3.008E-03 6.064E-03 7.578E-03 8.572E-03 8.998E-03 8.328E-03 9.023E-03 1.004E-02 1.107E-02
LNO Mean 1.679E-03 2.103E-02 5.276E-02 6.218E-02 7.448E-02 8.865E-02 9.384E-02 1.191E-01 1.124E-01 1.055E-01
Std 1.609E-04 8.680E-03 1.922E-02 2.044E-02 2.413E-02 2.766E-02 3.131E-02 3.564E-02 3.749E-02 3.263E-02
Transformer Mean 3.220E-02 2.431E-02 2.547E-02 2.623E-02 2.805E-02 2.969E-02 3.061E-02 3.453E-02 3.402E-02 3.262E-02
Std 3.003E-04 1.651E-04 8.204E-04 8.835E-04 1.904E-03 1.275E-03 2.416E-03 3.093E-03 3.925E-03 3.595E-03
Oformer-V Mean 6.087E-04 1.183E-03 2.366E-03 2.995E-03 3.703E-03 4.404E-03 4.291E-03 5.937E-03 6.083E-03 5.339E-03
Std 3.744E-05 5.219E-04 1.361E-03 1.684E-03 2.096E-03 2.574E-03 2.709E-03 3.598E-03 3.677E-03 3.385E-03
Oformer-G Mean 6.431E-04 2.932E-04 2.923E-04 2.897E-04 3.218E-04 3.316E-04 3.254E-04 3.868E-04 3.830E-04 3.618E-04
Std 7.398E-05 2.669E-05 1.716E-05 1.651E-05 1.122E-05 4.508E-05 3.668E-05 6.187E-05 5.957E-05 5.335E-05
Oformer-F Mean 6.310E-04 2.573E-04 2.341E-04 2.115E-04 2.250E-04 2.129E-04 2.065E-04 2.301E-04 2.286E-04 2.184E-04
Std 1.362E-04 3.097E-05 2.657E-05 3.138E-05 2.668E-05 3.435E-05 3.221E-05 3.509E-05 3.426E-05 3.533E-05
GNOT Mean 9.961E-04 5.128E-04 7.509E-04 8.636E-04 1.103E-03 1.196E-03 1.217E-03 1.640E-03 1.702E-03 1.541E-03
Std 9.282E-05 4.917E-05 1.052E-04 2.285E-04 3.337E-04 4.293E-04 3.227E-04 5.873E-04 6.608E-04 4.580E-04
Mamba Mean 1.086E-04 9.311E-05 1.034E-04 1.029E-04 1.461E-04 1.156E-04 1.111E-04 1.410E-04 1.626E-04 1.288E-04
Std 1.485E-05 2.302E-05 2.816E-05 2.869E-05 8.059E-05 3.776E-05 5.116E-05 4.718E-05 5.692E-05 6.369E-05